# AGAV-Rater: Adapting Large Multimodal Model for AI-Generated Audio-Visual Quality Assessment

### v1

@@ 5.5 Ablation Study
Effects of Pretraining Procedure. The “w/o pretraining” column in Tab. 4 shows AGAV-Rater’s performance without the pretraining step. Compared to the AGAVQA-MOS subset, the pretraining step significantly enhances AGAV-Rater’s performance on the smaller TTA and TTM datasets. This is because pretraining provides rich prior knowledge, which effectively compensates for the limited scale and diversity of these smaller datasets. On the AGAVQA-MOS subset, the pretraining step shows a more significant improvement in the overall A/V quality, compared to audio quality and A/V content consistency. We believe this is because AGAV-Rater relies on the scores of audio quality and A/V content consistency when predicting the overall A/V quality, and pretraining helps AGAV-Rater learn audio quality and A/V content consistency faster, reducing disturbances in learning overall A/V quality.
@@ 6 Conclusion
In this paper, we construct AGAVQA, the first AGAV quality assessment dataset, which labels AGAV quality in two ways: multi-dimensional score prediction and optimal AGAV selection. We propose a novel LMM-based AGAV quality assessment method, AGAV-Rater. AGAV-Rater demonstrates superior score prediction capabilities on the AGAVQA-MOS, TTA, and TTM datasets, and achieves the highest accuracy in identifying the optimal AGAV on the AGAVQA-Pair subset AGAV-Rater offer users a better audio-visual experience and enhance the quality of VTA method outputs.
Appendix A More Details of AGAVQA Construction.
@@ A.2 Detailed Information of VTA Methods
We utilize $16$ state-of-the-art VTA methods to construct the AGAVQA dataset, encompassing both open-source and closed-source methods. For open-source models, we rely on official repositories and utilize default weights to generate audio. For closed-source models, we utilize publicly accessible APIs provided by open platforms. For models without publicly available code, we collect public AGAVs showcased on their Github pages.
@@ A.2.5 Proprietory VTA Method
A.3 More Subjective Socres Analysis
In the AGAVQA-MOS subset, we collected MOSs across three dimensions, with the distribution of MOSs for each dimension shown in Fig. 4. MOSs are evenly distributed between $20$ and $80$ for all three dimensions. Additionally, as illustrated in Fig. 5, we compare the average MOSs of different VTA methods across various video audio sources. The audio generated by SVA consistently demonstrates superior quality across all source types. Most VTA methods show better audio quality in the instrument category. In the A/V content consistency dimension, most VTA methods struggle with the fantasy, cooking, people, and scenery categories, where it is challenging to generate audio that aligns with the content of these four types. In the overall A/V quality dimension, VTA methods underperform in the people and cooking categories.
Appendix B More Details of Experiments
B.1 Detailed Information of Compared Methods
@@
B.2 Cost Analysis of AGAV-Rater

### v2

@@ 5.5 Ablation Study
Effects of Pretraining Procedure. The “w/o Pretrain” column in Tab. 4 shows AGAV-Rater’s performance without the pretraining step. Compared to the AGAVQA-MOS subset, the pretraining step significantly enhances AGAV-Rater’s performance on the smaller TTA and TTM datasets. This is because pretraining provides rich prior knowledge, which effectively compensates for the limited scale and diversity of these smaller datasets. As shown in Tab. 1, VideoLLaMA2 exhibits relatively weak performance on the content consistency dimension for the TTM dataset. The pre-training procedure, through the music-text instruction-response pairs, enhances the music perception ability of AGAV-Rater, thereby improving its performance on the consistency dimension of the TTM dataset. On the AGAVQA-MOS subset, the pretraining step shows a more significant improvement in the overall A/V quality, compared to audio quality and A/V content consistency. We believe this is because AGAV-Rater relies on the scores of audio quality and A/V content consistency when predicting the overall A/V quality, and pretraining helps AGAV-Rater learn audio quality and A/V content consistency faster, reducing disturbances in learning overall A/V quality.
@@
Effects of the Base Models. We conduct ablation studies using GroundingGPT (Li et al., 2024) and OneLLM (Han et al., 2024) as base models on the AGAVQA-MOS subset. For OneLLM and GroundingGPT, we first load the default weights, and then fine-tune them using the official training code on the AGAVQA-MOS subset. To ensure fairness, we also add the quality regression module, directly regressing the LLM’s last hidden states to output three-dimensional numerical scores. As shown in Tab. 5, AGAV-Rater achieves the best performance. The main reason for this is that VideoLLaMA2 is designed for audio-video content understanding and pre-trained on more diverse audio-video datasets, making it more suitable for our quality assessment task. GroundingGPT focuses more on localization and visual understanding and is not designed or trained to understand continuous audio-video content. Its ability to comprehend video quality may be weaker. OneLLM is a general multimodal model that, while supporting audio and video processing, is not specifically optimized or enhanced for video and audio alignment. Its audio-related dataset only includes audio-text data, and OneLLM is more suited to text-vision or text-audio matching and understanding, rather than specific audio-video content.
@@ 6 Conclusion
In this paper, we construct AGAVQA-3k, the first AGAV quality assessment dataset, which labels AGAV quality in two ways: multi-dimensional score prediction and optimal AGAV selection. We propose a novel LMM-based AGAV quality assessment method, AGAV-Rater. AGAV-Rater demonstrates superior score prediction capabilities on the AGAVQA-MOS, TTA, and TTM datasets, and achieves the highest accuracy in identifying the optimal AGAV on the AGAVQA-Pair subset. AGAV-Rater enhances users a better audio-visual experience and enhance the quality of VTA method outputs.
Impact Statement
The quality of AI-generated audio and video must align with human preferences. Among existing models, AGAV-Rater achieves the highest consistency with human perceptual evaluations of AGAVs, indicating its potential for supervising and controlling the quality of AGAVs. We will continue to focus on and advance this research in the future.
Appendix A More Details of AGAVQA-3k Construction.
@@ A.2 Detailed Information of VTA Methods
We utilize $16$ state-of-the-art VTA methods to construct the AGAVQA-3k dataset, encompassing both open-source and closed-source methods. For open-source models, we rely on official repositories and utilize default weights to generate audio. For closed-source models, we utilize publicly accessible APIs provided by open platforms. For models without publicly available code, we collect public AGAVs showcased on their Github pages.
@@ A.2.5 Proprietory VTA Method
Figure 4: An example of the scoring interface. The audio-visual setup consists of two $1080$p monitors and a headphone. One monitor displays the scoring interface, while the other is used for subjects to watch the AGAV videos and listen to the audio through the headphones.
Appendix B More Details of Human Evaluation
B.1 Scoring Dimensions
During the subjective scoring process, participants are asked to evaluate the AGAVs from three dimensions: audio quality, A/V content consistency, and overall A/V quality. These three dimensions provide a comprehensive and detailed assessment of the AGAVs. Audio quality mainly focuses on the naturalness and clarity of the audio, minimizing the influence of the video content. Higher scores indicate that the audio is more natural, clear, and free of distortion, while lower scores reflect that distortions and noise in the sound degrade the listener’s auditory experience. The scoring range is from 00 to $5$, with scores accurate to one decimal place. The scoring criteria are as follows:
  • 4–5 (Excellent): The audio is natural, clear, and free of distortion.
  • 3–4 (Good): The audio is generally clear but contains slight distortion or noise.
  • 2–3 (Fair): The audio is somewhat muffled, with noticeable distortion or noise.
  • 1–2 (Poor): The audio is quite muffled, with severe distortion and noise that significantly affect the listening experience.
  • 0–1 (Bad): The audio is completely distorted.
A/V content consistency is independent of audio quality and primarily focuses on whether the sounds or music in the audio align with the video content. Higher scores indicate a strong correlation between the audio and the video content, while lower scores indicate a low correlation, with the audio lacking the sound elements intended to be conveyed by the video. The scoring criteria are as follows:
  • 4–5 (Excellent): The audio content is highly consistent with the video content.
  • 3–4 (Good): The audio content is generally consistent with the video content.
  • 2–3 (Fair): The audio content has limited correlation with the video content.
  • 1–2 (Poor): The audio content is mostly inconsistent with the video content.
  • 0–1 (Bad): The audio is largely distorted, and the audio content is completely inconsistent with the video content.
Overall A/V quality mainly focuses on the overall perceptual experience of the audio and video, including audio quality, A/V content consistency, and A/V temporal synchronization. Higher scores indicate clear and natural audio, with both content and timing highly consistent with the video. Lower scores indicate obvious audio distortion and low correlation with the video content. The scoring criteria are as follows:
  • 4–5 (Excellent): The overall quality of the audio and video is excellent.
  • 3–4 (Good): The overall quality of the audio and video is good, with slight distortion.
  • 2–3 (Fair): The overall quality of the audio and video is average, with noticeable distortion.
  • 1–2 (Poor): The overall audio and video quality is poor, with severe distortion or mostly inconsistent content.
  • 0–1 (Bad): The overall audio and video quality is very poor, with completely inconsistent content and severely distorted audio or video.
We developed a subjective evaluation guideline that includes the scoring criteria for the three dimensions mentioned above. Before each participant begins the subjective experiment, we guide them through reading this document to familiarize themselves with the scoring criteria for each dimension, thereby ensuring accuracy and consistency of the ratings across different participants.
B.2 Scoring Interface
The example of scoring interface is shown in Fig. 4. We designed the scoring interface using the Python Tkinter package. Before scoring, each subject is prompted to enter their username, which allows retrieval of their scoring progress and subsequent presentation of the scoring interface. The interface includes three continuous quality rating bars, three navigation buttons, and a display of the number of AGAVs that have been scored. Each rating bar is labeled with a 11115555 Likert scale, and participants can drag the slider to assign scores accurate to one decimal place. The navigation buttons, including “Prev”, “Repeat”, and “Next”, enable participants to modify the previous AGAV’s score, replay the current AGAV, and submit their score to proceed to the next AGAV, respectively, facilitating efficient scoring. All AGAVs are viewed at their original resolution without any scaling or cropping.
B.3 Evaluation Environment
The official testing phase was conducted in a controlled laboratory setting with normal indoor lighting and a quiet environment. Subjects were seated at a comfortable distance of approximately $60$ cm from the screen. We used a Redmi $23.8$-inch monitor with a resolution of $1920\times 1080$ and Sony WH-1000XM4 headphones to minimize distortions introduced by the viewing and listening devices during human evaluation.
B.4 Subject Selection
We invited subjects familiar with AVQA and AGAV to participate in on-site training sessions. Detailed explanations of the scoring criteria for each dimension were provided, along with additional AGAV samples for practice. Expert reviewers then evaluated the annotations and selected $15$ qualified subjects. Each subject rated all $3,088$ samples in the AGAVQA-MOS subset in a randomized order. To prevent fatigue, each subject was limited to rating a maximum of $60$ samples per day, completing the entire task in approximately two months. This protocol ensured the validity and robustness of each subject’s ratings.
@@ B.4 Subject Selection
Table 7: Standard deviation of subjective scores for each category in the AGAVQA-MOS subset.
Animal Water People Vehicle Object Scenery Sea Fantasy Fire Instrument Cooking
12.56 12.16 12.20 11.79 12.57 12.74 12.37 13.11 12.35 11.93 12.23
B.5 More Subjective Socres Analysis
In the AGAVQA-MOS subset, we collected MOSs across three dimensions. SRCC between audio quality and A/V content consistency is $0.6860$, indicating that the two dimensions are independent. SRCC between audio quality and A/V overall quality is $0.7876$, and between content consistency and overall quality is $0.7926$, suggesting that overall quality is influenced by both audio quality and content consistency.
The distribution of MOSs for each dimension shown in Fig. 5. MOSs are evenly distributed between $20$ and $80$ for all three dimensions. Additionally, as illustrated in Fig. 6, we compare the average MOSs of different VTA methods across various video audio sources. The audio generated by SVA consistently demonstrates superior quality across all source types. Most VTA methods show better audio quality in the instrument category. In the A/V content consistency dimension, most VTA methods struggle with the fantasy, cooking, people, and scenery categories, where it is challenging to generate audio that aligns with the content of these four types. In the overall A/V quality dimension, VTA methods underperform in the people and cooking categories.
The range of MOSs is $[0,100]$. We categorize the video content into $11$ main audio sources and then calculate the standard deviation of the overall quality scores among $15$ subjects. We compute the average standard deviation for each category, as shown in Tab. 7. The “Fantasy” category shows the highest standard deviation, as it represents unreal scenarios, leading to more diverse interpretations among subjects.
Appendix C More Details of Experiments
C.1 Detailed of Datasets
We conduct experiments on our proposed AGAVQA-3k dataset, as well as on the TTA and TTM datasets. The TTA dataset (Deshmukh et al., 2024) generates $500$ audio samples from $100$ prompts using $5$ text-to-audio generation methods. Subjects were then invited to rate the quality of the audio and its relevance to the provided description. Similarly, the TTM dataset (Deshmukh et al., 2024) generates $500$ music samples from $100$ prompts using $5$ text-to-music generation methods. Subjects were asked to rate the quality of the music and its relevance to the provided description.
The TTM dataset does not include an evaluation of music’s aesthetic quality. Since music aesthetic quality assessment is heavily influenced by personal preferences and the subject’s taste in different music styles, it is difficult to achieve objective evaluation. Therefore, in the AGAVQA-MOS, TTA, and TTM datasets, the audio quality dimension primarily focuses on the quality and realism of the audio.
C.2 Detailed of Loss Function
We first pre-train AGAV-Rater using text-defined levels. The model predicts a fixed-length text output of $1$ token, representing either “Excellent” or “Bad”. During pre-training, the cross-entropy loss is used as the objective function, which can be defined as:
$L_{CE}=-\frac{1}{N}\sum^{N}_{n=1}\sum^{C}_{c=1}y_{n,c}log(\hat{y}_{n,c}),$ (2)
where $N$ is the number of AGAVs in the batch, $C$ is the vocabulary size, $y_{n,c}$ is the ground-truth label for the $n$-th AGAV at vocabulary position $c$. $\hat{y}_{n,c}$ is the predicted probability output by the model at position $c$ for the $n$-th AGAV. For fine-tuning, we train AGAV-Rater using numerical scores and employ the Pearson Linear Correlation Coefficient (PLCC) loss as the objective. The PLCC loss is defined as:
$L=(1-\frac{\left<\widehat{s}-mean(\widehat{s}),s-mean(s)\right>}{\lVert\widehat{s}-mean(\widehat{s})\rVert_{2}\lVert s-mean(s)\rVert_{2}})/2,$ (3)
[…]
C.3 Detailed Information of Compared Methods
@@
C.4 Details of Audio Preprocessing
AGAV-Rater uses default parameters from VideoLLaMA2 for audio preprocessing. Assuming $T$ video frames are extracted, the steps are:
  1. 1.
    Divide audio into $T$ segments.
  2. 2.
    Concatenate all segments, then crop or zero-pad to a fixed length.
  3. 3.
    Transform into fbank spectrograms with $128$ frequency bins.
  4. 4.
    Use BEATs and an MLP block to extract features from the spectrograms.
  5. 5.
    Concatenate audio, video, and text features, and input them into the LLM.
Table 10: $230$ AIGC videos content distribution and AGAV-Rater filtering accuracy for each category. We collected these AIGC videos and then used ElevenLabs to generate AGAVs, which were subsequently filtered by AGAV-Rater to select high-quality AGAVs in Section 5.6.
Metric Animal Water People Vehicle Object Scenery Sea Fantasy Fire Instrument Cooking All
Video Number 41 15 20 22 31 19 20 11 15 24 12 230
AGAV-Rater filtering accuracy 0.78 0.87 0.85 0.77 0.81 0.79 0.80 0.82 0.80 0.83 0.75 0.80
C.5 More Details of Enhancing Elevenlabs Results via AGAV-Rater
In Section 5.6, We collected $230$ silent AIGC videos and used ElevenLabs to generate audios, aiming to verify that AGAV-Rater can assist ElevenLabs in selecting high-quality audio for users. We conducted a statistical analysis of the content distribution of these $230$ AIGC videos, with the results presented in Tab. 10. The analysis shows that the video content is quite diverse, enabling a comprehensive evaluation of AGAV-Rater’s performance across different types of video content. We further present the accuracy of AGAV-Rater in identifying higher-quality AGAVs across these categories. As shown in the Tab. 10, AGAV-Rater achieves over $75$% accuracy in each category.
C.6 Cost Analysis of AGAV-Rater